# Future Impact Decomposition in Request-level Recommendations

### v4

@@ Abstract.
${}^{\dagger}$ Work was done during an internship at Kuaishou Technology.
${}^{\ast}$ The first two authors contributed equally to this work.
${}^{\ddagger}$ Corresponding author
conference: XXX; XXX,XXX; XXX,XXX
@@ 1. Introduction
For practical reasons, the recommendation policy is typically designed to provide a list of items/contents rather than a single item as an action in order to efficiently process the frequent and continuous browsing requests (Ie et al., 2019). This design choice is often seen in scenarios like short video, news, and blog recommendation systems, where users may typically consume/browse a large number of contents in a short time. However, this list-wise recommendation mechanism is fundamentally inconsistent with the user’s item-wise behavior (we will use list-wise and request-level interchangeably in this paper). Intuitively, a user only pays attention to one item at a time and his/her actual “state” changes after the interaction with each item, but the system is assumed to infer a list, obtain a list-wise feedback, and then update the user’s state on each request.
In this paper, we emphasize this inconsistency in RL-based recommendation and formulate a request-level MDP with item-wise rewards. Specifically, the system can only observe how the list-wise state changes in the request-level MDP, but we assume that the ground-truth item-level state transition in the user’s viewpoint is not observable. A straightforward solution under the request-level MDP would simply optimize the cumulative list-wise reward using any standard RL approach (Mnih et al., 2016; Lillicrap et al., 2016), but this may induce information loss on the item characteristics for future reward estimation. As an intuitive example, Tom receives a list of movie recommendations from either sports or documentary categories and gives positive feedback to sports movies but no feedback to documentary ones. If using a list-level optimization, the system may not discover the item-level differences and Tom’s preference for sports. Fortunately, the item-wise user feedback throughout the interaction history is available, which augments the list-wise reward with a more detailed view. Yet, each items’ long-term value consists of both the immediate feedback and its future impact. And it is reasonable to believe that the causal effects of a certain item on the user’s future interactions(Gao et al., 2023) are not evenly distributed in the list of exposure. Following the previous example, recommending sports movies may satisfy Tom’s demands and gain more trust from him, so potentially increasing Tom’s engagement in the future. In contrast, recommending a documentary might be an unsatisfactory action as it could spoil the system’s impression to Tom and lead to a decrease in user engagement in the future. In general, we need an RL-based solution that can optimize the policy under the list-wise state transition, but also exploits both the item-level reward and item-wise future impact.
@@
To verify the aforementioned assumptions, we first propose an item-decomposed advantage actor-critic (ItemA2C) learning framework to distinguish items with different immediate feedback. The model optimizes the same cumulative reward in the request-level MDP as standard request-level A2C, but both the critic’s temporal difference loss and the actor’s advantage maximization loss are decomposed into an item-wise optimization. The empirical results demonstrate that this item-level paradigm is superior to its request-level counterpart. To further investigate how the item decomposition mechanism manipulates the optimization of the future impact of items, we design a re-weighting strategy where the evaluation of each item’s future value may vary with respect to its immediate reward. In other words, this strategy assumes that better immediate rewards indicate better future impact. The resulting framework theoretically still recovers the request-level actor-critic framework while providing separate item-wise long-term value estimation. Knowing that there may exist other valid choices in the re-weighting space, we generalize the heuristic design of weights into a weight model which is optimized through adversarial optimization.
@@ 2.1. List-wise Recommendation
The study on the list-wise recommendation problem mainly focuses on the items mutual influences (Cao et al., 2007; Xia et al., 2008; Burges, 2010; Ai et al., 2018), and early approaches only discuss this issue in a top-k recommendation scenario where the goal is to directly optimize a list-wise metric. Later research efforts also provide solutions to a more complicated sequential next-list recommendation scenario, where the list-wise MDP is introduced (Rendle et al., 2010; Ie et al., 2019; Liu et al., 2023a). As one of the most representative works, SlateQ (Ie et al., 2019) is a reinforcement learning method for recommender systems (RL4RS). It uses a DQN framework and proposes a decomposition technique that learns to distribute the list’s Q-value toward each individual item, and formulates the optimization problem of finding a slate with the maximum expected Q-value as a fractional linear program. Though SlateQ is effective in modeling the list-wise reward, its ”single-choice” assumption mismatches our setting where all items in the lists may have non-zero rewards. In general, we distinguish our problem setting from existing works as it emphasizes the existing item-wise feedback under the sequential list-wise recommendation scenario.
@@ 2.2. Reinforcement Learning for Recommendation
The RL-based recommendation system(RS) (Afsar et al., 2021; Sutton and Barto, 2018; Zhao et al., 2018a) adopts the MDP framework and aims to optimize the cumulative reward that reflects long-term user satisfaction. Although tabular-based (Mahmood and Ricci, 2007) methods can be used to optimize an evaluation table in a simple setting, they are limited to a fixed set of state-action pairs. Therefore, value-based methods (Pei et al., 2019; Taghipour et al., 2007; Zhao et al., 2021a, 2018b) and policy gradient methods (Chen et al., 2019b, a; Ge et al., 2021, 2022; Li et al., 2022; Xian et al., 2019) have been proposed to learn to evaluate and optimize the action policy based on the sampled long-term reward. The actor-critic (Zhao et al., 2018a; Xin et al., 2020; Zhao et al., 2020; Cai et al., 2023b) paradigm integrates these two methods by simultaneously learning an action evaluator and an action generator. Our method belongs to this later paradigm. Despite its effectiveness, reinforcement learning also faces several challenges when accommodating the recommender system, including but not limited to the exploration of combinatorial state/actions space (Dulac-Arnold et al., 2015; Ie et al., 2019; Liu et al., 2020, 2023b), unstable user behavior (Bai et al., 2019; Chen et al., 2021a), heterogeneous user feedback (Chen et al., 2021b; Cai et al., 2023b), and multi-task learning (Feng et al., 2018; Tang et al., 2020; Liu et al., 2023c). Among all existing works, we consider HAC (Liu et al., 2023b) as closest to our work since it also included item-wise components in an RL framework. In comparison, HAC aims to represent a list in a hyper-embedding space for end-to-end training, while our work focus on the item-level long-term effect modeling. Note that we also emphasize the importance of item-level future impact attribution which could compliment the HAC model.
@@ 4.1. Offline Experiments with Simulator
Table 1. Online simulation performance of all methods.The best performances in bold and second best in Underline.
type Model ML1M KuaiRand
Total Reward Depth Variance Total Reward Depth Variance
request-level DDPG 12.97 ±$\pm(2.95)$ 13.72 ±$\pm(2.62)$ 36.44±$\pm(2.40)$ 11.71 ±$\pm(0.98)$ 12.93 ±$\pm(0.83)$ 51.58 ±$\pm(7.12)$
A2C 16.88 ±$\pm(0.58)$ 17.32 ±$\pm(0.50)$ 20.25 ±$\pm(2.25)$ 10.01 ±$\pm(0.86)$ 11.56 ±$\pm(0.72)$ 71.82 ±$\pm(4.88)$
HAC 17.53 ±$\pm(0.13)$ 17.90 ±$\pm(0.12)$ 21.71 ±$\pm(3.76)$ 12.65±$\pm(0.26)$ 13.72±$\pm(0.22)$ 47.53±$\pm(4.26)$
item-level Supervision 12.57 ±$\pm(2.13)$ 13.71 ±$\pm(1.80)$ 51.14 ±$\pm(9.15)$ 9.78±$\pm(3.15)$ 11.32±$\pm(2.66)$ 52.30±$\pm(9.96)$
SlateQ 4.62 ±$\pm(0.19)$ 6.96 ±$\pm(0.16)$ 16.01 ±$\pm(0.82)$ 3.37 ±$\pm(0.20)$ 5.97 ±$\pm(0.17)$ 23.73 ±$\pm(1.11)$
ItemA2C($\alpha=0$) 17.58 ±$\pm(0.62)$ 17.94 ±$\pm(0.53)$ 18.72 ±$\pm(4.68)$ 13.45±$\pm(0.08)$ 14.41±$\pm(0.07)$ 48.30±$\pm(5.24)$
ItemA2C-W($\alpha=0.5$) 17.57 ±$\pm(0.54)$ 17.93 ±$\pm(0.46)$ 18.35 ±$\pm(2.95)$ 14.77 ±$\pm(0.41)$ 15.52±$\pm(0.35)$ 41.42±$\pm(4.58)$
ItemA2C-W($\alpha=1$) 17.80±$\pm(0.73)$ 18.12±$\pm(0.63)$ 16.46±$\pm(3.84)$ 15.48 ±$\pm(0.61)$ 16.13 ±$\pm(0.53)$ 35.26 ±$\pm(1.97)$
ItemA2C-M 17.94±$\pm(0.46)$ 18.23±$\pm(0.40)$ 15.88±$\pm(2.40)$ 16.03±$\pm(0.53)$ 16.59±$\pm(0.45)$ 28.11±$\pm(2.62)$
4.1.1. Datasets and Online Simulator
4.1.2. Evaluation Protocol
With the online simulator ready, we can apply pseudo-online training for RL models, and evaluate its performance in the last 100 episode steps (minimum 320 user episodes). We observed that the majority of RL-based techniques achieved convergence and stabilization within 30,000 iterations. For long-term evaluation metrics, we report the total reward (without discount) and depth of user sessions/episodes. Higher values for these metrics indicate superior performance. To evaluate the stability of the learned policies, we included a reward variance metric that estimates inconsistency within 10 episode steps. Thus, lower variance demonstrates greater stability.
4.1.3. Baselines
  • A2C(section 3.2): The request-level Actor-Critic method.
We present the details of the specifications in Appendix D.
4.1.4. Model Alternatives
We denote our item-decomposition framework with equal weights as itemA2C, the reward-based re-weighting strategy as itemA2C-W, and the extended framework that uses weight model with adversarial learning is noted as itemA2C-M(model). For reproduction of our empirical study, we provide implementation and training details in our released source code  [^1] https://anonymous.4open.science/r/ItemDecomposition.
    (a)  Performance in ML1M         (b)  Performance in KuaiRand
Figure 3. Learning curves of all methods.
4.1.5. Main Results
For each experiment across all models, we run online training and evaluation for five rounds with different random seeds and report the average results in Table 1. We can see that most RL-based methods (DDPG, A2C, HAC, and ItemA2C) can improve the recommendation result over supervised learning, indicating the superiority of optimizing the cumulative reward over immediate reward. The SlateQ method applies item decomposition but is not well suited for our problem setting since it decomposes the request-level total reward without using the item-wise reward, which results in sub-optimal performance. Among the baselines, the best request-level method is HAC in terms of the total reward and depth metrics. Including the itemwise supervision might be the reason why HAC surpasses other baselines, but its lack of item-wise decomposition of future impact is an inferior design which is reflected in the empirical results compared to our itemA2C alternatives. In comparison, ItemA2C-M with weight model achieves the best overall performance and significantly (with $p<0.05$) improves the reward by 27% and the depth by 20% in KuaiRand1K and marginally (though not statistically significant) improves the reward by 2.3% and depths by 1.8% in ML1M. Currently, we believe the difference between the improvement made on ML1M and KuaiRand is related to the consistency of user behaviors (more consistent interactions in the short video platform than that in the movie) and the number of candidate items (3706 items in ML1M and 11643 items in KuaiRand). These data characteristics make the item selection process easier for the baseline models on ML1M dataset to approach the optimal performance. Though not the focus of this paper, further investigation on the notion of user behavior consistency could be a valuable extension to all RL methods.
4.1.6. Strategic vs. Model-based Reweighting
4.1.7. Ablation
Effect of list size: Additionally, the list size is set to $K=6$ by default but it is reasonable to believe that the choice of list size also affects the recommendation result. In our empirical study on both datasets, ItemA2C consistently achieves better performances than the best baseline HAC and all other methods for $K>2$. Moreover, the superiority of ItemA2C-M over ItemA2C-W is also verified for different settings of $K$. We present a detailed analysis of different list size choices in Appendix B.
4.2. Live Experiments
Table 2. Online performances of item-level A2C
Method Watch Time Like Follow Collect Comment
[…]
4.2.1. Experimental Setup
4.2.2. Results Analysis
Figure 7. Total reward under different $\alpha$. Each experiment was run for five times with different random seeds, and the results include mean, median, quartiles, and min-max.
5. Conclusion and Discussion
In this paper, we present the challenge of inconsistent viewpoints between request-level recommendation policy and item-level user behaviors. To address this challenge, we formulated a practical list-wise MDP formulation incorporating item-wise reward and proposed an item-wise decomposition method under the actor-critic learning framework. Based on the proposed method, we show that one should not only utilize the item-wise reward for the optimization of each immediate step but also distinguish each items future impact through the value function decomposition. The proposed method introduces no extra learnable parameters than A2C and is well-suited for item-wise parallel training. We additionally proposed a model-based method with adversarial learning that automatically learns to decompose the future impact and proves that the reward-based strategy is reasonable but could be further improved. And following the same notion in our model-based approach, future research on new ways to represent the future impact of each item may also give insights into user behavior modeling, improving the recommendation systems and user satisfaction.
Appendix A Analysis on Item Decomposition
list size ML1M
HAC itemA2C($\alpha=0$) itemA2C-W($\alpha=0.5$) ItemA2C-M
Reward Depth Reward Depth Reward Depth Reward Depth
6 17.60±$\pm(0.33)$ 17.96±$\pm(0.28)$ 17.75±$\pm(0.58)$ 18.08±$\pm(0.50)$ 17.57±$\pm(0.54)$ 17.93±$\pm(0.46)$ 17.94±$\pm(0.47)$ 18.23±$\pm(0.40)$
list size KuaiRand
HAC itemA2C($\alpha=0$) itemA2C-W($\alpha=0.5$) ItemA2C-M
Reward Depth Reward Depth Reward Depth Reward Depth
6 12.65±$\pm(0.26)$ 13.72±$\pm(0.22)$ 13.49±$\pm(0.41)$ 14.45±$\pm(0.35)$ 14.77±$\pm(0.61)$ 15.52±$\pm(0.52)$ 16.03±$\pm(0.53)$ 16.59±$\pm(0.45)$
Table 3. Performance for different list size.
A.1. An Intuitive Example
In this section, we use some examples to show how different item decomposition strategies may influence the learning process. Considering the relation between request-level reward and item-level reward:
(10) $R(s_{t},a_{t})=\sum_{i_{t,k}\in a_{t}}r_{t,k}$
Suppose we observe item-wise click-or-not reward as
$[\frac{1}{2},0,0,\frac{1}{2},0,0]$ if $\ \ \ \ \text{if }\alpha=1.0$
Appendix B Effects of List Size
The main experiment in our paper set the list size to $K=6$ to mirror the real scenario presented in the online system. However, it is worthwhile to see how the decomposition mechanism behaves under different list sizes. We conducted experiments with $K\in\{1,2,4,6,8\}$ which changes both the MDP of the RL model but also the user environment. We test both request-level A2C and the ItemA2C (with equal weights and re-weighting $\alpha=0.5$), and run each experiment for five rounds with different random seeds and report the average values in Table 3. The performance of all methods gradually deteriorated as the list size increases. This is a direct indicator of the existence of the inconsistent view between request-level system MDP and item-level user MDP, and the most accurate view aligns with the item-wise user view. The proposed ItemA2C-WA achieves the best result across different list size settings.
@@ Appendix B Effects of List Size
Appendix C Full main results
Table 4 presents the extended results of the main experiments, including batch-wise average, minimum, and maximum value for both the reward and depth metric. Notably, the ItemA2C-WA achieved the best performance across all of these metrics in both ML1M and KuaiRand1K.
Appendix D The Details of all Methods

### v5

@@ Abstract.
Work was done during an internship at Kuaishou Technology.
The first two authors contributed equally to this work.
Corresponding author
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 25–29, 2024; Barcelona, Spainbooktitle: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’24), August 25–29, 2024, Barcelona, Spaindoi: 10.1145/3637528.3671506isbn: 979-8-4007-0490-1/24/08ccs: Information systems Recommender systemsccs: Theory of computation Reinforcement learning
@@ 1. Introduction
For practical reasons, the recommendation policy is typically designed to provide a list of items/contents rather than a single item as an action in order to efficiently process the frequent and continuous browsing requests (Ie et al., 2019). This design choice is often seen in scenarios like short video, news, and blog recommendation systems, where users may typically consume/browse a large number of contents in a short time. However, this list-wise recommendation mechanism is fundamentally inconsistent with the user’s item-wise behavior. Intuitively, a user only pays attention to one item at a time and his/her actual “state” changes after the interaction with each item, but the system is assumed to infer a list, obtain a list-wise feedback, and then update the user’s state on each request.
In this paper, we emphasize this inconsistency in RL-based recommendation and formulate a request-level MDP with item-wise rewards. Specifically, the system can only observe how the state changes in the request-level MDP, but we assume that the ground-truth item-level state transition in the user’s viewpoint is not observable. A straightforward solution under the request-level MDP would simply optimize the cumulative list-wise reward using any standard RL approach (Mnih et al., 2016; Lillicrap et al., 2016), but this may induce information loss on the item characteristics for future reward estimation. As an intuitive example, Tom receives a list of movie recommendations from either sports or documentary categories and gives positive feedback to sports movies but no feedback to documentary ones. If using a list-level optimization, the system may not discover the item-level differences and Tom’s preference for sports. Fortunately, the item-wise user feedback throughout the interaction history is available, which augments the list-wise reward with a more detailed view. Yet, each items’ long-term value consists of both the immediate feedback and its future impact. And it is reasonable to believe that the causal effects of a certain item on the user’s future interactions(Gao et al., 2023) are not evenly distributed in the list of exposure. Following the previous example, recommending sports movies may satisfy Tom’s demands and gain more trust from him, so potentially increasing Tom’s engagement in the future. In contrast, recommending a documentary might be an unsatisfactory action as it could spoil the system’s impression to Tom and lead to a decrease in user engagement in the future. In general, we need an RL-based solution that can optimize the policy under the list-wise state transition, but also exploits both the item-level reward and item-wise future impact.
@@
To verify the aforementioned assumptions, we first propose an item-decomposed advantage actor-critic (ItemA2C) learning framework to distinguish items with different immediate feedback. The model optimizes the same cumulative reward in the request-level MDP as standard request-level A2C, but both the critic’s temporal difference loss and the actor’s advantage maximization loss are decomposed into an item-wise optimization. The empirical results demonstrate that this item-level paradigm is superior to its request-level counterpart. To further investigate how the item decomposition mechanism manipulates the optimization of the future impact of items, we design a re-weighting strategy where the evaluation of each item’s future value may vary with respect to its immediate reward. In other words, this strategy assumes that better immediate rewards indicate better future impact. The resulting framework theoretically recovers the request-level actor-critic framework while providing separate item-wise long-term value estimation. Knowing that there may exist other valid choices in the re-weighting space, we generalize the heuristic design of weights into a weight model which is optimized through adversarial optimization.
@@ 2.1. List-wise Recommendation
For recommendation scenarios that take a list of items as an action, they can generally categorized in to one-step list recommendation and sequential list-wise recommendation problems. The first study mainly focuses on the items mutual influences (Cao et al., 2007; Xia et al., 2008; Burges, 2010; Ai et al., 2018), and early approaches only discuss this issue in a top-k recommendation scenario where the goal is to directly optimize a list-wise metric. The latter research efforts provide solutions to a more complicated sequential next-list recommendation scenario, where the list-wise MDP is introduced (Rendle et al., 2010; Ie et al., 2019; Liu et al., 2023a). As one of the most representative works, SlateQ (Ie et al., 2019) is a reinforcement learning method for recommender systems (RL4RS). It uses a DQN framework and proposes a decomposition technique that learns to distribute the list’s Q-value toward each individual item, and formulates the optimization problem of finding a slate with the maximum expected Q-value as a fractional linear program. Though SlateQ is effective in modeling the list-wise reward, its ”single-choice” assumption mismatches our setting where all items in the lists may have non-zero rewards. In general, we distinguish our problem setting from existing works as it emphasizes the existing item-wise feedback under the sequential list-wise recommendation scenario.
@@ 2.2. Reinforcement Learning for Recommendation
The RL-based recommendation system(RS) (Afsar et al., 2021; Sutton and Barto, 2018; Zhao et al., 2018a) adopts the MDP framework and aims to optimize the cumulative reward that reflects long-term user satisfaction. Although tabular-based (Mahmood and Ricci, 2007) methods can be used to optimize an evaluation table in a simple setting, they are limited to a fixed set of state-action pairs. Therefore, value-based methods (Pei et al., 2019; Taghipour et al., 2007; Zhao et al., 2021a, 2018b) and policy gradient methods (Chen et al., 2019b, a; Ge et al., 2021, 2022; Li et al., 2022; Xian et al., 2019) have been proposed to learn to evaluate and optimize the action policy based on the sampled long-term reward. The actor-critic (Zhao et al., 2018a; Xin et al., 2020; Zhao et al., 2020; Cai et al., 2023b) paradigm integrates these two methods by simultaneously learning an action evaluator and an action generator. Our method belongs to this later paradigm. Despite its effectiveness, reinforcement learning also faces several challenges when accommodating the recommender system, including but not limited to the exploration of combinatorial state/actions space (Dulac-Arnold et al., 2015; Ie et al., 2019; Liu et al., 2020, 2023b), unstable user behavior (Bai et al., 2019; Chen et al., 2021a), heterogeneous user feedback (Chen et al., 2021b; Cai et al., 2023b), and multi-task learning (Meng et al., 2023; Liu et al., 2024, 2023c). Among all existing works, we consider HAC (Liu et al., 2023b) as closest to our work since it also included item-wise components in an RL framework. In comparison, HAC aims to represent a list in a hyper-embedding space for end-to-end training, while our work focus on the item-level long-term effect modeling. Note that we also emphasize the importance of item-level future impact attribution which could compliment the HAC model.
@@ 4.1. Offline Experiments with Simulator
4.1.1. Datasets and Online Simulator
4.1.2. Evaluation Protocol
With the online simulator ready, we can apply pseudo-online training for RL models, and evaluate its performance in the last 100 episode steps (minimum 320 user episodes). We observed that the majority of RL-based techniques achieved convergence and stabilization within 30,000 iterations. For long-term evaluation metrics, we report the total reward (without discount) and depth of user sessions/episodes. Higher values for these metrics indicate superior performance.
4.1.3. Baselines
  • A2C(section 3.2): The request-level Actor-Critic method.
We present the details of the specifications in Appendix C.
Table 2. Performance for different list size.
list size ML1M
HAC itemA2C($\alpha=0$) itemA2C-W($\alpha=0.5$) ItemA2C-M
Reward Depth Reward Depth Reward Depth Reward Depth
16 15.65±$\pm(0.74)$ 15.96±$\pm(0.69)$ 16.12±$\pm(0.96)$ 16.39±$\pm(0.91)$ 16.23±$\pm(0.91)$ 16.43±$\pm(0.90)$ 16.27±$\pm(0.99)$ 16.51±$\pm(0.93)$
32 14.72±$\pm(0.82)$ 14.93±$\pm(0.78)$ 14.79±$\pm(1.02)$ 15.00±$\pm(0.98)$ 14.58±$\pm(0.73)$ 14.68±$\pm(0.67)$ 14.85±$\pm(1.13)$ 15.05±$\pm(1.09)$
list size KuaiRand
HAC itemA2C($\alpha=0$) itemA2C-W($\alpha=0.5$) ItemA2C-M
Reward Depth Reward Depth Reward Depth Reward Depth
16 11.51±$\pm(0.35)$ 11.94±$\pm(0.30)$ 11.28±$\pm(0.57)$ 11.93±$\pm(0.53)$ 12.74±$\pm(0.26)$ 13.37±$\pm(0.24)$ 13.73±$\pm(0.77)$ 14.18±$\pm(0.71)$
32 10.37±$\pm(0.22)$ 10.71±$\pm(0.19)$ 9.71±$\pm(0.64)$ 10.11±$\pm(0.61)$ 11.29±$\pm(0.36)$ 11.63±$\pm(0.33)$ 11.39±$\pm(0.41)$ 11.71±$\pm(0.39)$
4.1.4. Model Alternatives
We denote our item-decomposition framework with equal weights as itemA2C, the reward-based re-weighting strategy as itemA2C-W, and the extended framework that uses weight model with adversarial learning is noted as itemA2C-M(model). For reproduction of our empirical study, we provide implementation and training details as well as the hyperparameters of best results in our grid search space in our released source code  [^1] https://github.com/wangxiaobei565/ItemDecomposition.
    (a)  Performance in ML1M         (b)  Performance in KuaiRand
Figure 3. Learning curves of all methods.
4.1.5. Main Results
For each experiment across all models, we run online training and evaluation for five rounds with different random seeds and report the average results in Table 1. We can see that most RL-based methods (DDPG, A2C, HAC, and ItemA2C) can improve the recommendation result over supervised learning, indicating the superiority of optimizing the cumulative reward over immediate reward. The SlateQ method applies item decomposition but is not well suited for our problem setting since it decomposes the request-level total reward without using the item-wise reward, which results in sub-optimal performance. Among the baselines, the best request-level method is HAC in terms of the total reward and depth metrics. Including the itemwise supervision might be the reason why HAC surpasses other baselines, but its lack of item-wise decomposition of future impact is an inferior design which is reflected in the empirical results compared to our itemA2C alternatives. In comparison, ItemA2C-M with weight model achieves the best overall performance and significantly (with $p<0.05$) improves the reward by 27% and the depth by 20% in KuaiRand1K and marginally (though not statistically significant) improves the reward by 2.3% and depths by 1.8% in ML1M. Currently, we believe the difference between the improvement made on ML1M and KuaiRand is related to the consistency of user behaviors (more consistent interactions in the short video platform than that in the movie) and the number of candidate items (3706 items in ML1M and 11643 items in KuaiRand). These data characteristics make the item selection process easier for the baseline models on ML1M dataset to approach the optimal performance.
4.1.6. Strategic vs. Model-based Reweighting
4.1.7. Ablation
Effect of list size: The main experiment in our paper set the list size to $K=6$ to mirror the real scenario presented in the online system. but it is reasonable to believe that the choice of list size also affects the recommendation result. We conducted experiments with $K\in\{1,2,4,6,8,16,32\}$ which changes both the MDP of the RL model but also the user environment. We test both request-level A2C and item-level ItemA2C, ItemA2C-W, ItemA2C-M and run each experiment for five rounds with different random seeds and report the average values in Table 2. The performance of all methods gradually deteriorated as the list size increases. This is a direct indicator of the existence of the inconsistent view between request-level system MDP and item-level user MDP, and the most accurate view aligns with the item-wise user view. The proposed ItemA2C-M achieves the best result across different list size settings. In our empirical study on both datasets, ItemA2C consistently achieves better performances than the best baseline HAC and all other methods for $K>2$. Moreover, the superiority of ItemA2C-M over ItemA2C-W is also verified for different settings of $K$.
4.2. Live Experiments
Table 3. Online performances of item-level A2C and all results are statistically significant.
Method Watch Time Like Follow Collect Comment
[…]
4.2.1. Experimental Setup
4.2.2. Results Analysis
Figure 7. Total reward under different $\alpha$ and the results include mean, median, quartiles, and min-max.
5. Conclusion and Discussion
In this paper, we present the challenge of inconsistent viewpoints between request-level recommendation policy and item-level user behaviors. To address this challenge, we formulated a practical list-wise MDP formulation incorporating item-wise reward and proposed an item-wise decomposition method under the actor-critic learning framework. Based on the proposed method, we show that one should not only utilize the item-wise reward for the optimization of each immediate user feedback but also distinguish each item’s future impact through the value function decomposition. The proposed method introduces no extra learnable parameters than A2C and is well-suited for item-wise parallel training. We additionally proposed a model-based method with adversarial learning that automatically learns to decompose the future impact and proves that the reward-based strategy is reasonable but could be further improved. Note that the model-based re-weighting solution may introduce extra modeling and training cost but no extra computation during inference. In terms of the data collection, one has to construct the data stream for the request-level MDP with item-wise feedback to implement the solution. Fortunately, this data is usually easy to collect in most industrial recommender systems. As an extension to our model-based approach, future research on new ways to represent the future impact of each item may also give insights into user behavior modeling, improving the recommendation systems and user satisfaction.
Additionally, our design does not restrict the reward settings to click-based reward design as in the offline experiment or linear combination of multi-response as in the online A/B test, future research on more complicated non-linear reward design (e.g. intra-list diversity) may further improve recommender performance. During deployment, our approach is well-suited for the last ranking stage or the reranking stage where user feedback are provided for all the output items, and some extra efforts might needed to accommodate previous ranking stages, since the candidate pool size is much larger and the output size are not necessarily K. In the RL’s viewpoint, our proposed method naturally extends the A2C framework, and we believe it is also worthwhile to investigate valid designs for future impact decomposition under DDPG or HAC.
Appendix A Analysis on Item Decomposition
A.1. An Intuitive Example
In this section, we use some examples to show how different item decomposition strategies may influence the learning process. Considering the relation between request-level and item-level reward:
(10) $R(s_{t},a_{t})=\sum_{i_{t,k}\in a_{t}}r_{t,k}$
Suppose we observe item-wise click-or-not reward as
$[\frac{1}{2},0,0,\frac{1}{2},0,0]$ if $\ \ \ \ \text{if }\alpha=1.0$
Appendix B Decomposition for critic learning
Table 4. Component analysis of ItemA2C where $\alpha=0.5$ in re-weighting strategy.
Dataset Performance ItemA2C ItemA2C-W(A) ItemA2C-W(C) ItemA2C-M(A)
ML1M Reward 17.58±$\pm(0.62)$ 17.57±$\pm(0.54)$ 17.57±$\pm(0.54)$ 17.94±$\pm(0.47)$
Depth 17.94±$\pm(0.53)$ 17.93±$\pm(0.46)$ 17.92±$\pm(0.46)$ 18.24±$\pm(0.40)$
KuaiRand Reward 13.45±$\pm(0.08)$ 14.77±$\pm(0.54)$ 13.41±$\pm(0.40)$ 16.03±$\pm(0.53)$
Depth 14.41±$\pm(0.07)$ 15.52±$\pm(0.35)$ 14.38±$\pm(0.45)$ 16.59±$\pm(0.45)$
In our method, the critic is following the TD learning objective of request level optimization where can be easily decomposed by future impact. We provide reward performance of this alternative in the ”itema2c-W(C)” column of the following table. It is not even surpassing itemA2C (without reweighting). We believe that this may be related to the fact that we deploy the actor instead of the critic during recommendation. Theoretically, it is possible to decompose the critic learning into item-level, but the items in the same state will eventually aggregate their results under request-level MDP.
Appendix C The Details of all Methods